OpenStudy (jacobbirddog):

Write an equation in slope-intercept form for the line passing through (–1, 10) and (5, 8).

OpenStudy (anonymous):

the equation for slope is y = mx + b. this is slope intercept form

OpenStudy (anonymous):

m is the slope, and b is the y intercept.

OpenStudy (jacobbirddog):

how would i write that?

OpenStudy (anonymous):

subtract the x1 and y 1 then x2 and y2

OpenStudy (anonymous):

that is your slope.

OpenStudy (abb0t):

Find the slope first: \(\sf slope=m=\frac{\Delta y}{\Delta x}\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

ABBOT LET ME HANDLE THIS ONE

OpenStudy (anonymous):

lol jk take it away !!

OpenStudy (abb0t):

Then, pick ANY point of the two, and plug it into: \(\sf \color{Red}{(y-y_1)=m(x-x_1)}\) and rearrange it into the form: \(\sf \color{purple}{y=mx+b}\)

OpenStudy (jacobbirddog):

Im not good with algebra

OpenStudy (abb0t):

You are pluggig it in for \(\sf \color{RED}{y_1 , x_1}\) NOT for y

OpenStudy (abb0t):

I already told you how to solve. I am not doing the work for you. Try it first. I told you step by step.